# On the quantitative lower bounds for solutions to the Boltzmann equation in nonconvex domains

### v3

@@ 6. Lower bound for the non-cutoff case
In this chapter, we consider lower bounds for the solution of the Boltzmann equation in the non-cutoff case, that is, $\nu\geq 0$. Here, the domain $\Omega$ is not necessarily convex. As in the cutoff case, we decompose the collision operator, but in a different way:
(6.1) $\begin{split}Q(h_{1},h_{2})(v)&=\int_{\mathbb{R}^{3}\times\mathbb{S}^{2}}B(|v-v_{*}|,\cos{\theta})[h_{2}(v^{\prime})h_{1}(v^{\prime}_{*})-h_{2}(v)h_{1}(v_{*})]dv_{*}d\sigma\\ &=\int_{\mathbb{R}^{3}\times\mathbb{S}^{2}}B(|v-v_{*}|,\cos{\theta})[h_{1}(v^{\prime}_{*})(h_{2}(v^{\prime})-h_{2}(v))]dv_{*}d\sigma\\ &\hskip 42.67912pt-h_{2}(v)\int_{\mathbb{R}^{3}\times\mathbb{S}^{2}}B(|v-v_{*}|,\cos{\theta})(h_{1}(v_{*})-h_{1}(v^{\prime}_{*}))dv_{*}d\sigma\\ &=:Q^{1}_{b}(h_{1},h_{2})(v)-Q^{2}_{b}(h_{1},h_{2})(v),\end{split}$
@@
$Q(h_{1},h_{2})=Q^{+}_{\epsilon}(h_{1},h_{2})-Q^{-}_{\epsilon}(h_{1},h_{2})+Q^{1}_{\epsilon}(h_{1},h_{2})+Q^{2}_{\epsilon}(h_{1},h_{2}),$
@@
$Q^{1}_{\epsilon}[h_{1},h_{2}](v):=h_{2}(v)\int_{\mathbb{R}^{3}\times\mathbb{S}^{2}}\Phi(|v-v_{*}|)b(\cos{\theta})\mathbf{1}_{\theta<\epsilon}[h_{1}(v_{*})-h_{1}(v^{\prime}_{*})]dv_{*}d\sigma.$
@@
$b^{CO}_{\epsilon}(\cos{\theta}):=\mathbf{1}_{\theta\geq\epsilon}b(\cos{\theta}),\,b^{NCO}_{\epsilon}(\cos{\theta}):=\mathbf{1}_{\theta<\epsilon}b(\cos{\theta}),$
$n_{b^{CO}_{\epsilon}}:=\int_{\mathbb{S}^{2}}b^{CO}_{\epsilon}(\cos{\theta})d\sigma,\,m_{b^{NCO}_{\epsilon}}:=\int_{\mathbb{S}^{2}}b^{NCO}_{\epsilon}(\cos{\theta})(1-\cos{\theta})d\sigma,$
@@
$S(h)(v):=\int_{\mathbb{R}^{3}\times\mathbb{S}^{2}}\Phi(|v-v_{*}|)b(\cos{\theta})[h(v_{*})-h(v^{\prime}_{*})]dv_{*}d\sigma.$
$S_{\epsilon}(h)(v):=\int_{\mathbb{R}^{3}\times\mathbb{S}^{2}}\Phi(|v-v_{*}|)b(\cos{\theta})\mathbf{1}_{\theta<\epsilon}[h(v_{*})-h(v^{\prime}_{*})]dv_{*}d\sigma.$
@@
We consider $\Omega\subset\mathbb{R}^{3}$ satisfying Assumption A, a kernel $B$ satisfying Assumption B with $\nu\geq 0$. Given a non negative continuous function $f_{0}$ on $\overline{\Omega}\times\mathbb{R}^{3}$, we call a non negative continuous function $f$ on $0,T)\times(\overline{\Omega}\times\mathbb{R}^{3})$ with $|f(t,x,v)|\leq C(1+|v|)^{r}$ for some constant $C>0$ and $r>3$ for any $0<t\leq T$, $(x,v)\in\overline{\Omega}\times\mathbb{R}^{3}$ a ”continuous mild” solution of (1.1)–(1.3) with initial data $f_{0}$ if there exists a number $0<\epsilon_{0}<\frac{\pi}{4}$ such that for any $0<\epsilon<\epsilon_{0}$, $(t,x,v)\in[0,T)\times\Omega\times\mathbb{R}^{3}$, we have
(6.3) $\begin{split}&f(t,x,v)=f_{0}(X_{0,t}(x,v),v)\exp\bigg[-\int_{0}^{t}(L_{\epsilon}+S_{\epsilon})[f(s,X_{s,t}(x,v),\cdot)](v)ds\bigg]\\ +&\int_{0}^{t}\exp\left(-\int_{s}^{t}(L_{\epsilon}+S_{\epsilon})[f(s^{\prime},X_{s^{\prime},t}(x,v),\cdot)](v)ds^{\prime}\right)\\ &(Q^{+}_{\epsilon}+Q_{\epsilon}^{1})[f(s,X_{s,t}(x,v),\cdot),f(s,X_{s,t}(x,v),\cdot)](v)ds,\end{split}$
when $t\leq t_{\partial}:=\sup\{t|x-vs\in\overline{\Omega},\ \forall s\in[0,t]\}$, and
(6.4) $\begin{split}f(t,x,v)&=\alpha f(t-t_{\partial},X_{t-t_{\partial},t}(x,v),R(X_{t-t_{\partial},t}(x,v),v))\\ &\exp\left[-\int_{t-t_{\partial}}^{t}(L_{\epsilon}+S_{\epsilon})[f(s,X_{s,t}(x,v),\cdot)](v)ds\right]\\ &+(1-\alpha)\left(\int_{w\cdot n(X_{t-t_{\partial},t}(x,v))>0}f(t,X_{t-t_{\partial},t}(x,v),w)(w\cdot n(X_{t-t_{\partial},t}(x,v)))dw\right)\\ &\frac{1}{2\pi T_{B}^{2}}e^{-\frac{|v|^{2}}{2T_{B}}}\exp\left[-\int_{t-t_{\partial}}^{t}(L_{\epsilon}+S_{\epsilon})[f(s,X_{s,t}(x,v),\cdot)](v)ds\right]\\ +&\int_{t-t_{\partial}}^{t}\exp\left(-\int_{s}^{t}(L_{\epsilon}+S_{\epsilon})[f(s^{\prime},X_{s^{\prime},t}(x,v),\cdot)](v)ds^{\prime}\right)\\ &(Q^{+}_{\epsilon}+Q_{\epsilon}^{1})[f(s,X_{s,t}(x,v),\cdot),f(s,X_{s,t}(x,v),\cdot)](v)ds\end{split}$
when $t\geq t_{\partial}$.
Before introducing the main result in the non cutoff case, we introduce the following constants (here $\widetilde{\gamma}:=(2+\gamma)^{+}$):
@@
We consider $\Omega\subset\mathbb{R}^{3}$ satisfying Assumption A, a kernel $B$ satisfying Assumption B with $\nu\geq 0$, $\alpha\in[0,1]$, and a non negative function $f_{0}$ which is continuous on $\overline{\Omega}\times\mathbb{R}^{3}$. We also consider a continuous mild solution $f(t,x,v)$ of (1.1)–(1.3) on $0,T)\times\overline{\Omega}\times\mathbb{R}^{3}$, with $T>0$ (and initial condition $f_{0}$) satisfying the following property:
@@
    $E_{f}<\infty$ if $\gamma\geq 0$ and $\max\{E_{f},L_{f}^{p_{\gamma}}\}<\infty$, where $p_{\gamma}>\frac{3}{3+\gamma}>0$, if $3<\gamma<0$;
@@
[…]
    and $E^{\prime}_{f}:=\sup\limits_{(t,x)\in[0,T)\times\Omega}\int_{\mathbb{R}^{3}}|v|^{(2+\gamma)^{+}}f(t,x,v)dv<\infty$.
Then,the following lower bound holds: For any $\tau\in(0,T)$, $K>2\frac{\log{(2+\frac{2\nu}{2-\nu})}}{\log{2}}$, there exist $C_{1}>0$, $C_{2}>0$ which depend on $C_{\Phi},c_{\Phi},\gamma,b_{0},\nu,E_{f}$ $W_{f}$$E^{\prime}_{f}$, $M$ (and $L_{f,p}$ if $\gamma<0$ ), $\tau$, $K$, and the modulus of continuity of $f$, such that
(6.9) $f(t,x,v)\geq C_{1}e^{-C_{2}|v|^{K}},\ \forall t\in[\tau,T),\ \forall(x,v)\in\overline{\Omega}\times\mathbb{R}^{3}.$
@@
Consider $g$ a measurable function on $\mathbb{R}^{3}$, and assume that the collision operator satisfies Assumption B with $0\leq\nu<2$. Then,there exists $C_{g}^{S}>0$ which depends only on $m_{b}$, $C_{\Phi}$ and $e_{g}$ (and $l_{g}^{p}$, where $p>\frac{3}{3+\gamma}$, if $\gamma<0$) such that
(6.10) $|S[g](v)|\leq C_{g}^{S}\langle v\rangle^{\gamma^{+}}.$
Here we define $m_{b}:=\int_{\mathbb{S}^{2}}b(\cos{\theta})(1-\cos{\theta})d\sigma$.
@@
(6.11) $|Q^{1}_{b}\left(h_{1},h_{2}\right)|\leq Cm_{b}c_{\Phi}\|h_{2}\|_{L^{1}_{\widetilde{\gamma}}}\|h_{1}\|_{W^{2,\infty}},$
@@
We consider $\Omega\subset\mathbb{R}^{3}$ satisfying Assumption A, a kernel $B$ satisfying Assumption B with $0\leq\nu<2$, $\alpha\in[0,1]$, and a mild continuous solution $f(t,x,v)$ of (1.1)–(1.3). Suppose that $A>0$, $\Delta_{1},\ \Delta_{2}>0$, $(\tau,x^{\prime},v^{\prime})\in[0,T)\times\Omega\times\mathbb{R}^{3}$ are such that $B(x^{\prime},\Delta_{2})\subset\Omega$ and
@@
where the numbers $\{r_{n}(\Delta_{2})\}_{n=0}^{\infty}\in\mathbb{R}$, $\{t_{n}(\tau,t,\Delta_{2},|v^{\prime}|)\}_{n=0}^{\infty}\in\mathbb{R}$ and $\{\alpha^{NC}_{n}(\tau,t,\Delta_{2},A,|v^{\prime}|)\}_{n=1}^{\infty}\in\mathbb{R}$ are defined as below:
@@
(6.14) $t_{n}(\tau,t,\Delta_{2},|v^{\prime}|):=\max\{\tau,t-\frac{\Delta_{2}}{2^{n+1}(2r_{n}(\Delta_{2})+|v^{\prime}|)}\},$
@@
(6.16) $\begin{split}&\alpha^{NC}_{n+1}(\tau,t,\Delta_{2},A,|v^{\prime}|):=Cr_{n}^{3+\gamma}(\Delta_{2})\int_{t_{n}(\tau,t,\Delta_{2},|v^{\prime}|)}^{t}e^{-(t-s)C<2r_{n}(\Delta_{2})+|v^{\prime}|>^{\gamma^{+}}}\alpha^{NC}_{n}(\tau,s,\Delta_{2},A,|v^{\prime}|)^{2}ds\\ &=Cr_{n}^{3+\gamma}(\Delta_{2})\int_{0}^{t-t_{n}(\tau,t,\Delta_{2},|v^{\prime}|)}e^{-uC<2r_{n}(\Delta_{2})+|v^{\prime}|>^{\gamma^{+}}}\alpha^{NC}_{n}(\tau,t-u,\Delta_{2},A,|v^{\prime}|)^{2}du,\end{split}$
where $C$ is a constant which depends on $\gamma,\ \nu,\ b_{0},\ n_{b}$, $m_{b}$ $C_{\Phi}$ and $E_{f}$, $W_{f}$, $E^{\prime}_{f}$ MM(and $L_{f,p}$, where $p>\frac{3}{3+\gamma}$, if $\gamma<0$).
@@
We prove the proposition by induction on $n$. The case $n=0$ is exactly the assumption. Assume that the statement of Proposition 6.2 holds for $n=k$. Given $t\in[\tau,\tau+\Delta_{1}]$, $x\in B(x^{\prime},\frac{\Delta_{2}}{2^{k+1}})$, $v\in B(0,|v^{\prime}|+2r_{k}(\Delta_{2}))$, we first notice that when $s\in[\max\{\tau,t-\frac{\Delta_{2}}{2^{k+1}(2r_{k}+|v^{\prime}|)}\},t]$, we have
@@
which implies that $X_{s,t}(x,v)\in B(x^{\prime},\frac{\Delta_{2}}{2^{k}})\subset\Omega$ and we deduce by induction hypothesis that
(6.17) $\forall v\in\mathbb{R}^{3},\ f(t,x,v)\geq\alpha^{NC}_{k}(\tau,t,\Delta_{2},A,|v^{\prime}|)\mathbf{1}_{B(v^{\prime},r_{k}(\Delta_{2}))}(v).$
We consider the second term on the right-hand side of (6.3) to obtain the following lower bound which holds for $0<\epsilon<\epsilon_{0}$ ($\epsilon_{0}$ is mentioned in Definition 6.1):
(6.18) $\begin{split}&f(t,x,v)\\ &\geq\int_{\tau}^{t}\exp\left(-\int_{s}^{t}(L_{\epsilon}+S_{\epsilon})[f(s^{\prime},X_{s^{\prime},t}(x,v),\cdot)](v))ds^{\prime}\right)\\ &(Q^{+}_{\epsilon}-|Q_{\epsilon}^{1}|)[f(s,X_{s,t}(x,v),\cdot),f(s,X_{s,t}(x,v),\cdot)](v)ds.\end{split}$
Now, by Lemma 3.1 and Lemma 6.1, we have
(6.19) $|(L_{\epsilon}+S_{\epsilon})[f(s^{\prime},X_{s^{\prime},t}(x,v),\cdot)](v)|\leq(C_{L,1}m_{b^{CO}_{\epsilon}}+C_{L,2}m_{b^{NCO}_{\epsilon}})\langle|v^{\prime}|+2r_{k}(\Delta_{2})\rangle^{\gamma^{+}},$
where the constants $C_{L,1},C_{L,2}$ depend only on $n_{b}$, $m_{b}$, $C_{\Phi}$ and $E_{f}$ (and $L_{f,p}$ if $\gamma<0$).
Hence, the following estimate holds:
(6.20) $\begin{split}&f(t,x,v)\\ &\geq\int_{\tau}^{t}\exp\left(-(t-s)(C_{L,1}m_{b_{CO}}+C_{L,2}m_{b_{NCO}})\langle|v^{\prime}|+2r_{k}(\Delta_{2})\rangle^{\gamma^{+}}\right)\\ &(Q^{+}_{\epsilon}-|Q_{\epsilon}^{1}|)[f(s,X_{s,t}(x,v),\cdot),f(s,X_{s,t}(x,v),\cdot)](v)ds,\end{split}$
for all $\frac{\pi}{4}>\epsilon>0$.
Next, we obtain as in the cutoff case that
(6.21) $\begin{split}&Q_{\epsilon}^{+}[f(s,X_{s,t}(x,v),\cdot),f(s,X_{s,t}(x,v),\cdot)](v)\\ &\\ &\geq C_{Q,1}\alpha^{NC}_{k}(\tau,s,\Delta_{2},A,|v^{\prime}|)^{2}l_{b^{CO}_{\epsilon}}c_{\Phi}(r_{k}(\Delta_{2}))^{3+\gamma}\xi^{\frac{1}{2}}\mathbf{1}_{B(v^{\prime},r_{k}(\Delta_{2})\sqrt{2}(1-\xi))}\\ &=C_{Q,1}\alpha^{NC}_{k}(\tau,s,\Delta_{2},A,|v^{\prime}|)^{2}l_{b^{CO}_{\epsilon}}c_{\Phi}(r_{k}(\Delta_{2}))^{3+\gamma}\xi^{\frac{1}{2}},\end{split}$

### v4

@@ 6. Lower bound for the non-cutoff case
In this chapter, we investigate lower bounds for the solution of the Boltzmann equation in the non-cutoff case, that is, $\nu\geq 0$. Here, the domain $\Omega$ is not necessarily convex. In contrast to the cutoff case studied in Chapter 2–4, the non-cutoff setting requires a refined decomposition of the collision operator to handle the angular singularity:
(6.1) $\begin{split}Q[h_{1},h_{2}](v)&=\int_{\mathbb{R}^{3}\times\mathbb{S}^{2}}B(|v-v_{*}|,\cos{\theta})[h_{2}(v^{\prime})h_{1}(v^{\prime}_{*})-h_{2}(v)h_{1}(v_{*})]\,dv_{*}\,d\sigma\\ &=\int_{\mathbb{R}^{3}\times\mathbb{S}^{2}}B(|v-v_{*}|,\cos{\theta})[h_{1}(v^{\prime}_{*})(h_{2}(v^{\prime})-h_{2}(v))]\,dv_{*}\,d\sigma\\ &\hskip 42.67912pt-h_{2}(v)\int_{\mathbb{R}^{3}\times\mathbb{S}^{2}}B(|v-v_{*}|,\cos{\theta})(h_{1}(v_{*})-h_{1}(v^{\prime}_{*}))\,dv_{*}\,d\sigma\\ &=:Q^{1}_{b}[h_{1},h_{2}](v)-Q^{2}_{b}[h_{1},h_{2}](v),\end{split}$
@@
$Q[h_{1},h_{2}](v)=Q^{+}_{\epsilon}[h_{1},h_{2}](v)-Q^{-}_{\epsilon}[h_{1},h_{2}](v)+Q^{1}_{\epsilon}[h_{1},h_{2}](v)+Q^{2}_{\epsilon}[h_{1},h_{2}](v),$
@@
$Q^{2}_{\epsilon}[h_{1},h_{2}](v):=h_{2}(v)\int_{\mathbb{R}^{3}\times\mathbb{S}^{2}}\Phi(|v-v_{*}|)b(\cos{\theta})\mathbf{1}_{\theta<\epsilon}[h_{1}(v_{*})-h_{1}(v^{\prime}_{*})]\,dv_{*}\,d\sigma.$
@@
(6.3) $b^{CO}_{\epsilon}(\cos{\theta}):=\mathbf{1}_{\theta\geq\epsilon}b(\cos{\theta}),\,b^{NCO}_{\epsilon}(\cos{\theta}):=\mathbf{1}_{\theta<\epsilon}b(\cos{\theta}),$
(6.4) $m_{b}:=\int_{\mathbb{S}^{2}}b(\cos{\theta})(1-\cos{\theta})\,d\sigma,$
(6.5) $n_{b^{CO}_{\epsilon}}:=\int_{\mathbb{S}^{2}}b^{CO}_{\epsilon}(\cos{\theta})\,d\sigma,\,m_{b^{NCO}_{\epsilon}}:=\int_{\mathbb{S}^{2}}b^{NCO}_{\epsilon}(\cos{\theta})(1-\cos{\theta})\,d\sigma,$
@@
$S[h](v):=\int_{\mathbb{R}^{3}\times\mathbb{S}^{2}}\Phi(|v-v_{*}|)b(\cos{\theta})[h(v_{*})-h(v^{\prime}_{*})]\,dv_{*}\,d\sigma,$
$S_{\epsilon}[h](v):=\int_{\mathbb{R}^{3}\times\mathbb{S}^{2}}\Phi(|v-v_{*}|)b(\cos{\theta})\mathbf{1}_{\theta<\epsilon}[h(v_{*})-h(v^{\prime}_{*})]\,dv_{*}\,d\sigma.$
@@
Suppose that $\Omega\subset\mathbb{R}^{3}$ satisfies Assumption A and the kernel $B$ satisfies Assumption B with $\nu\geq 0$. Given a non-negative continuous function $f_{0}$ on $\overline{\Omega}\times\mathbb{R}^{3}$, we call a non-negative continuous function $f$ defined on $0,T)\times(\overline{\Omega}\times\mathbb{R}^{3})$ with $|f(t,x,v)|\leq C(1+|v|)^{-r}$ for some constant $C>0$ and $r>3$ for any $0<t\leq T$, $(x,v)\in\overline{\Omega}\times\mathbb{R}^{3}$ a ”continuous mild” solution to (1.1)–(1.3) with initial data $f_{0}$ if $f$ is continuous on $\Gamma_{conti}$ and there exists a number $0<\epsilon_{0}<\frac{\pi}{4}$ such that for any $0<\epsilon<\epsilon_{0}$, $(t,x,v)\in[0,T)\times\Omega\times\mathbb{R}^{3}$, we have
(6.6) $\begin{split}f(t,x,v)&=f_{0}(X_{0,t}(x,v),v)\exp\bigg[-\int_{0}^{t}(L_{\epsilon}+S_{\epsilon})[f(s,X_{s,t}(x,v),\cdot)](v)\,ds\bigg]\\ &+\int_{0}^{t}\exp\left(-\int_{s}^{t}(L_{\epsilon}+S_{\epsilon})[f(s^{\prime},X_{s^{\prime},t}(x,v),\cdot)](v)\,ds^{\prime}\right)\\ &(Q^{+}_{\epsilon}+Q_{\epsilon}^{1})[f(s,X_{s,t}(x,v),\cdot),f(s,X_{s,t}(x,v),\cdot)](v)\,ds,\end{split}$
when $t\leq t_{\partial}(x,v)$, and
(6.7) $\begin{split}f(t,x,v)&=\alpha f(t-t_{\partial}(x,v),X_{t-t_{\partial}(x,v),t}(x,v),R(X_{t-t_{\partial}(x,v),t}(x,v),v))\\ &\exp\left[-\int_{t-t_{\partial}(x,v)}^{t}(L_{\epsilon}+S_{\epsilon})[f(s,X_{s,t}(x,v),\cdot)](v)\,ds\right]\\ &+(1-\alpha)\left(\int_{w\cdot n(X_{t-t_{\partial}(x,v),t}(x,v))>0}f(t,X_{t-t_{\partial}(x,v),t}(x,v),w)(w\cdot n(X_{t-t_{\partial}(x,v),t}(x,v)))dw\right)\\ &\frac{1}{2\pi T_{B}^{2}}e^{-\frac{|v|^{2}}{2T_{B}}}\exp\left[-\int_{t-t_{\partial}(x,v)}^{t}(L_{\epsilon}+S_{\epsilon})[f(s,X_{s,t}(x,v),\cdot)](v)\,ds\right]\\ &+\int_{t-t_{\partial}(x,v)}^{t}\exp\left(-\int_{s}^{t}(L_{\epsilon}+S_{\epsilon})[f(s^{\prime},X_{s^{\prime},t}(x,v),\cdot)](v)\,ds^{\prime}\right)\\ &(Q^{+}_{\epsilon}+Q_{\epsilon}^{2})[f(s,X_{s,t}(x,v),\cdot),f(s,X_{s,t}(x,v),\cdot)](v)\,ds\end{split}$
when $t\geq t_{\partial}(x,v)$. Here, we recall the definition of $t_{\partial}(x,v)$ from Definition 1.1.
Before introducing the main result in the non-cutoff case, we introduce the following constants (here $\widetilde{\gamma}:=(2+\gamma)^{+}$):
@@
Suppose that $\Omega\subset\mathbb{R}^{3}$ satisfies Assumption A and that the kernel $B$ satisfies Assumption B with $\nu\geq 0$, $\alpha\in[0,1]$. We consider a non-negative function $f_{0}$ that is continuous on $\overline{\Omega}\times\mathbb{R}^{3}$. Let $f(t,x,v)$ be a continuous mild solution to (1.1)–(1.3) on $0,T)\times\overline{\Omega}\times\mathbb{R}^{3}$, with $T>0$ (and the initial condition $f_{0}$) which satisfies the following properties:
@@
    $E_{f}<\infty$, where $p_{\gamma}>\frac{3}{3+\gamma}>0$, if $3<\gamma<0$;
@@
    $W_{f}<\infty$ and $E^{\prime}_{f}<\infty$.
Then, the following lower bound holds: There exists $0<\tau_{0}\leq T$ such that for any $\tau\in(0,\tau_{0})$ and $K>2\frac{\log{(2+\frac{2\nu}{2-\nu})}}{\log{2}}$, there exist $\Delta_{\tau_{0}}>0$, $\rho>0$ and $\theta>0$ depending on $\tau_{0}$, $C_{\Phi},c_{\Phi},\gamma,b_{0},\nu,E_{f}$ $W_{f}$$E^{\prime}_{f}$, $M$ (and $L_{f,p}$ if $\gamma<0$ ), $\tau$, $K$, and the modulus of continuity of $f_{0}$, such that
(6.12) $f(t,x,v)\geq\frac{\rho}{(2\pi\theta)^{\frac{3}{2}}}e^{-\frac{|v|^{K}}{2\theta}},\ \forall\,t\in[\tau,\Delta_{\tau_{0}}),\ \forall\,(x,v)\in\overline{\Omega}\times\mathbb{R}^{3}.$
@@
Consider $g$ a measurable function on $\mathbb{R}^{3}$, and assume that the collision operator satisfies Assumption B with $0\leq\nu<2$. Then, there exists $C_{g}^{S}>0$ which depends only on $m_{b}$, $C_{\Phi}$ and $e_{g}$ (and $l_{g,p}$, where $p>\frac{3}{3+\gamma}$, if $\gamma<0$) such that
(6.13) $|S[g](v)|\leq C_{g}^{S}\left\langle v\right\rangle^{\gamma^{+}}.$
Here we recall the definition of $e_{g}$ in (1.16), $l_{g,p}$ in (1.17) and define $m_{b}:=\int_{\mathbb{S}^{2}}b(\cos{\theta})(1-\cos{\theta})\,d\sigma$.
@@
(6.14) $|Q^{1}_{b}\left(h_{1},h_{2}\right)|\leq Cm_{b}c_{\Phi}\|h_{2}\|_{L^{1}_{\widetilde{\gamma}}}\|h_{1}\|_{W^{2,\infty}}.$
@@
Suppose that $\Omega\subset\mathbb{R}^{3}$ satisfies Assumption A and the kernel $B$ satisfies Assumption B with $0\leq\nu<2$, $\alpha\in[0,1]$. We consider a mild continuous solution $f(t,x,v)$ of (1.1)–(1.3). Given $A>0$, $0<\Delta_{1}\leq 1,$ $\Delta_{2}<1$, and $(\tau,x^{\prime},v^{\prime})\in[0,T)\times\Omega\times\mathbb{R}^{3}$ such that $B(x^{\prime},\Delta_{2})\subset\Omega$ and
@@
Then, there exist $\{\epsilon_{i}>0\}_{i=0}^{\infty}$ such that for $n\in\mathbb{N}\cup\{0\}$, $t\in[\tau,\tau+\Delta_{1}]$, $x\in B(x^{\prime},\frac{\Delta_{2}}{2^{n}})$,
@@
where the numbers $\{r_{n}(\Delta_{2})\}_{n=0}^{\infty}\in\mathbb{R}$, $\{t_{n}(\tau,t,\Delta_{2},|v^{\prime}|)\}_{n=0}^{\infty}\in\mathbb{R}$ and $\{\alpha^{NC}_{n}(\tau,t,\Delta_{2},A,|v^{\prime}|)\}_{n=0}^{\infty}\in\mathbb{R}$ are defined as follows:
@@
(6.17) $t^{NC}_{n}(\tau,t,\Delta_{2},|v^{\prime}|):=\max\left\{\tau,t-\frac{\Delta_{2}}{2^{n+1}(2r_{n}(\Delta_{2})+|v^{\prime}|)}\right\},$
@@
(6.19) $\begin{split}&\alpha^{NC}_{n+1}(\tau,t,\Delta_{2},A,|v^{\prime}|)\\ \ :=&\frac{1}{4}\int_{t^{NC}_{n}(\tau,t,\Delta_{2},|v^{\prime}|)}^{t}\exp\left(-(t-s)(C_{L,1}n_{b^{CO}_{\epsilon_{n}}}+C_{L,2}m_{b^{NCO}_{\epsilon_{n}}})\left\langle|v^{\prime}|+2r_{n}(\Delta_{2})\right\rangle^{\gamma^{+}}\right)\\ &\hskip 14.22636ptC_{Q,1}\alpha^{NC}_{n}(\tau,s,\Delta_{2},A,|v^{\prime}|)^{2}l_{b^{CO}_{\epsilon_{n}}}c_{\Phi}(r_{n}(\Delta_{2}))^{3+\gamma}\,ds.\end{split}$
where $C_{L,1}$, $C_{L,2}$ and $C_{Q,1}$ are constants which depend on $\gamma,\ \nu,\ b_{0},\ n_{b}$, $m_{b}$, $C_{\Phi}$, $E_{f}$, $W_{f}$, $E^{\prime}_{f}$, MM(and $L_{f,p}$, where $p>\frac{3}{3+\gamma}$, if $\gamma<0$).
@@
We prove the proposition by induction on $n$. The case $n=0$ is exactly the assumption. Assume that the statement of Proposition 6.2 holds for $n=k$. Given $t\in[\tau,\tau+\Delta_{1}]$, $x\in B\left(x^{\prime},\frac{\Delta_{2}}{2^{k+1}}\right)$, $v\in B\left(v^{\prime},r_{k+1}(\Delta_{2})\right)\subset B\left(0,|v^{\prime}|+2r_{k}(\Delta_{2})\right)$, we consider the second term on the right-hand side of (6.6) to obtain the following lower bound which holds for $0<\epsilon<\epsilon_{0}$ ($\epsilon_{0}$ is mentioned in Definition 6.1):
(6.20) $\begin{split}&f(t,x,v)\\ =&f(t^{NC}_{k}(\tau,t,\Delta_{2},|v^{\prime}|),X_{t^{NC}_{k}(\tau,t,\Delta_{2},|v^{\prime}|),t}(x,v),v)\\ &\exp\bigg[-\int_{t^{NC}_{k}(\tau,t,\Delta_{2},|v^{\prime}|)}^{t}(L_{\epsilon}+S_{\epsilon})[f(s,X_{s,t}(x,v),\cdot)](v)\,ds\bigg]\\ &+\int_{t^{NC}_{k}(\tau,t,\Delta_{2},|v^{\prime}|)}^{t}\exp\left(-\int_{s}^{t}(L_{\epsilon}+S_{\epsilon})[f(s^{\prime},X_{s^{\prime},t}(x,v),\cdot)](v)\,ds^{\prime}\right)\\ &(Q^{+}_{\epsilon}+Q_{\epsilon}^{1})[f(s,X_{s,t}(x,v),\cdot),f(s,X_{s,t}(x,v),\cdot)](v)\,ds,\\ \geq&\int_{t^{NC}_{k}(\tau,t,\Delta_{2},|v^{\prime}|)}^{t}\exp\left(-\int_{s}^{t}(L_{\epsilon}+S_{\epsilon})[f(s^{\prime},X_{s^{\prime},t}(x,v),\cdot)](v))\,ds^{\prime}\right)\\ &(Q^{+}_{\epsilon}-|Q_{\epsilon}^{1}|)[f(s,X_{s,t}(x,v),\cdot),f(s,X_{s,t}(x,v),\cdot)](v)\,ds.\end{split}$
Now, by Lemma 3.1 and Lemma 6.1, we have
(6.21) $|(L_{\epsilon}+S_{\epsilon})[f(s^{\prime},X_{s^{\prime},t}(x,v),\cdot)](v)|\leq(C_{L,1}n_{b^{CO}_{\epsilon}}+C_{L,2}m_{b^{NCO}_{\epsilon}})\left\langle|v^{\prime}|+2r_{k}(\Delta_{2})\right\rangle^{\gamma^{+}}$
for some constants $C_{L,1},C_{L,2}$ depending only on $n_{b}$, $m_{b}$, $C_{\Phi}$ and $E_{f}$ (and $L_{f,p}$ if $\gamma<0$).
Hence, the following estimate holds:
(6.22) $\begin{split}&f(t,x,v)\\ \geq&\int_{t^{NC}_{k}(\tau,t,\Delta_{2},|v^{\prime}|)}^{t}\exp\left(-(t-s)(C_{L,1}n_{b^{CO}_{\epsilon}}+C_{L,2}m_{b^{NCO}_{\epsilon}})\left\langle|v^{\prime}|+2r_{k}(\Delta_{2})\right\rangle^{\gamma^{+}}\right)\\ &(Q^{+}_{\epsilon}-|Q_{\epsilon}^{1}|)[f(s,X_{s,t}(x,v),\cdot),f(s,X_{s,t}(x,v),\cdot)](v)\,ds,\end{split}$
for all $\frac{\pi}{4}>\epsilon>0$.
We notice that when $s\in\left[t^{NC}_{k}(\tau,t,\Delta_{2},|v^{\prime}|),t\right]$, we have
@@
which implies that $X_{s,t}(x,v)\in B\left(x^{\prime},\frac{\Delta_{2}}{2^{k}}\right)\subset\Omega$ and we deduce by induction hypothesis that
(6.23) $\forall v\in\mathbb{R}^{3},\ f(s,X_{s,t}(x,v),v)\geq\alpha^{NC}_{k}(\tau,t,\Delta_{2},A,|v^{\prime}|)\mathbf{1}_{B(v^{\prime},r_{k}(\Delta_{2}))}(v).$
Next, we obtain from Lemma 3.2 by setting $\xi=\frac{1}{4}$: for any $v\in B\left(v^{\prime},r_{k+1}(\Delta_{2})\right)$, we have
(6.24) $\begin{split}&Q_{\epsilon}^{+}[f(s,X_{s,t}(x,v),\cdot),f(s,X_{s,t}(x,v),\cdot)](v)\\ \geq&C_{Q,1}\alpha^{NC}_{k}(\tau,s,\Delta_{2},A,|v^{\prime}|)^{2}l_{b^{CO}_{\epsilon}}c_{\Phi}(r_{k}(\Delta_{2}))^{3+\gamma}\frac{1}{2}\mathbf{1}_{B\left(v^{\prime},\frac{3\sqrt{2}}{4}r_{k}(\Delta_{2})\right)}\\ =&C_{Q,1}\alpha^{NC}_{k}(\tau,s,\Delta_{2},A,|v^{\prime}|)^{2}l_{b^{CO}_{\epsilon}}c_{\Phi}(r_{k}(\Delta_{2}))^{3+\gamma}\frac{1}{2},\end{split}$